40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 12266     Entry time: Thu Jul 7 12:44:52 2016
Author: varun 
Type: Update 
Category: CDS 
Subject: DAFI update 

Attached is a diagram, showing the entire (planned) signal flow of the DAF model. Some thoughts on the implementation after discussion with eric:

1) Since the LSC control signals and ASC signals are running on the c1lsc FE at the same rate as DAFI (16kHz), it would be wise to start from these.

   Current implementation: has a matrix at the end of the LSC PD signals, which selects one of the PD signals and outputs it to the DAFI via IPC communication.

    Proposed Changes: Add another matrix at the end of the LSC PD signals, to give to the second stereo output. Similarly, add two matrices each at the end of the LSC control signals and     the ASC signals. Each matrix must select one of the signals and output it to the DAF via IPC.

2) The PEM running on the c1sus FE system will have to be brought to DAFI in a similar fashon. However, since c1sus runs at 2kHz, there is a possibility of imaging while the signal is    transfered to the DAFI. This could be taken care of by an anti imaging filter, or inserting zeros between two samples coming at to the 16 kHz system from the 2kHz system and then low-passing it to remove the aliased parts. (similar to upsampling)

3) For the SUS control signals, input can be given from a matrix prepared for each optic seperately.

Attachment 1: DAFI.pdf  219 kB  | Hide | Hide all
DAFI.pdf
ELOG V3.1.3-